OpenStudy (anonymous):

Simplify. Please be sure to show all of your work. -4[-7 – (-2 + -8)]

OpenStudy (anonymous):

-4[-7 – (-2 + -8)] = -4[-7 – (-2 -8)] -4[-7 – (-10)] = -4[-7 +10] = -4[3] = -12

OpenStudy (anonymous):

thanks for the answer i see what i did wrong i did not do -4[-7 – (-10)]